# Quantum Generative Diffusion Model: A Fully Quantum-Mechanical Model for Generating Quantum State Ensemble

### v2

@@ III-B Forward Process
In the natural world, diffusion is governed by non-equilibrium thermodynamics, driving systems toward equilibrium [55]. As a system undergoes evolution, its entropy increases until it reaches maximum entropy at equilibrium, indicating that there is no further evolution. The classical diffusion model draws inspiration from this diffusion mechanism [28]. In quantum information theory, the depolarizing channel is considered as a “worst-case scenario” [56]. Under its influence, a quantum state gradually loses its original information, eventually transforming into a completely mixed state. This state is characterized by the highest entropy, representing the pinnacle of randomness in quantum systems. At this point, our uncertainty about the system has reached its maximum.
@@
where $d=2^{N}$ denotes the dimension of Hilbert space and $\alpha_{t}\in[0,1]$ is a timestep-dependent scalar. Eq. (1) indicates that the input quantum state $\rho_{t-1}$ remains unchanged with a probability of $\alpha_{t}$, while with a probability of $1-\alpha_{t}$, it is replaced by the completely mixed state $\mathbb{I}/d$. This indicates the loss of original information in the quantum system. Therefore, the depolarizing channel is non-unitary because information is lost when quantum states pass through it, making the process irreversible. By setting an appropriate sequence of $\{\alpha_{t}\}$, a series of depolarizing channel sequences acting on the quantum state $\rho_{0}$ will eventually yield the output $\rho_{T}\approx\mathbb{I}/d$.
III-B1 Efficient Quantum State Transition to Arbitrary Timestep
@@
where $\bar{\alpha}_{t}=\prod\limits_{i=1}^{t}\alpha_{i}$.
@@
$=\left(1-\bar{\alpha}_{t}\right)\mathbb{I}/d+\bar{\alpha}_{t}\rho_{0}.$
@@
III-B2 Noise Schedule
The noise parameter ${\alpha_{t}}$ determines how fast $\rho_{0}$ approaches $\rho_{T}$. The choice of ${\alpha_{t}}$ must ensure that $\bar{\alpha}_{T}\to 0$ and ${\alpha_{t}}\in\left[{0,1}\right]$ for any $t$. We use the cosine noise schedule from the classical diffusion model paper[31] to select $\alpha_{t}$:
@@ III-B2 Noise Schedule
with $s$ being a small offset hyperparameter. We note that there are other noise schedules, such as the linear noise schedule [29] and the learned noise schedule [57]. Exploring the impacts of other noise schedules on QGDM’s performance, or designing specific noise schedules for QGDM, is an interesting research direction, which we leave for future work.
III-C Backward Process
@@
The objective of the backward process in QGDM is to restore $\rho_{0}$ from $\rho_{T}$. We should address two considerations in the design of the backward process.
@@
Firstly, since the denoising process is essentially the inverse of the non-unitary diffusion process, it must also be non-unitary to effectively denoise. Secondly, in principle, we need to train $T$ distinct denoising processes to deal with $T$ different denoising demands if we model the denoising process in a form $f_{\boldsymbol{\Theta}_{t}}(\rho_{t})$. To reduce the number of trainable parameters in QGDM, we make these processes share the same parameters by using the timestep $t$ as an input. This integration is crucial for shaping the function $f_{\boldsymbol{\Theta}}(\rho_{t},t)$, as it allows the model to effectively utilize temporal information to accurately estimate $\rho_{t-1}$ while maintaining a low number of parameters.
@@
Figure 3: The backward process framework, $f_{\boldsymbol{\Theta}}(\rho_{t},t)$, is composed of a timestep embedding circuit $\mathcal{T}(\boldsymbol{\omega},t^{\prime})$ and a denoising circuit $U(\boldsymbol{\theta})$, where $\boldsymbol{\Theta}=\{\boldsymbol{\omega},\boldsymbol{\theta}\}$.
@@
To endow the denoising process with the aforementioned two characteristics, we achieve non-unitary properties by tracing out part of the subsystem. Additionally, we use a parameterized circuit to embed temporal information into the quantum system. Our design for the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ is illustrated in Fig. 3. $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ comprises two modules: a timestep embedding circuit $\mathcal{T}(\boldsymbol{\omega},t^{\prime})$ parameterized by $\boldsymbol{\omega}$ and $t^{\prime}=t\pi/T$, and a denoising circuit $U(\boldsymbol{\theta})$ parameterized by $\boldsymbol{\theta}$, thus trainable parameters $\boldsymbol{\Theta}=\{\boldsymbol{\omega},\boldsymbol{\theta}\}$. A timestep embedding circuit $\mathcal{T}(\boldsymbol{\omega},t^{\prime})$, operating on an auxiliary register with $N_{\text{time}}$ qubits, is utilized to produce state $\tau_{t}$, which carries temporal information. Subsequently, the composite system $\tau_{t}\otimes\rho_{t}$ is fed into the denoising circuit $U({\boldsymbol{\theta}})$. Afterward, all qubits are regrouped into two registers: register A with $N$ qubits and register B with $N_{\text{time}}$ qubits. We trace out the register B to get the predicted denoised quantum state $\tilde{\rho}_{t-1}$.
III-C1 Timestep Embedding Circuit
@@
A critical consideration when designing the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ is how to encode the temporal information $t$ into a quantum state. This is a classical-to-quantum encoding challenge, where methods such as qubit encoding [58, 59, 60] or amplitude encoding [61] could be used to embed $t$ into a quantum system. However, the qubit encoding method only maps the timestep sequence uniformly onto a ring on the Bloch sphere, which lacks expressiveness and flexibility. The amplitude encoding method requires a number of gates that is exponential in the number of qubits [62]. To enhance the expressiveness and flexibility of timestep embedding while maintaining an appropriate gate complexity, in this work, we utilize the quantum embedding method [63] to obtain a timestep embedding state $\tau_{t}$ with $N_{\text{time}}$ qubits:
@@
$\tau_{t}=\mathcal{T}(\boldsymbol{\omega},t^{\prime})\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\text{time}}}\mathcal{T}^{\dagger}(\boldsymbol{\omega},t^{\prime}),$ (6)
where scalar $t^{\prime}=t\pi/T$ is the mapping of timestep $t$ into the range [0, $\pi$], and $\boldsymbol{\omega}$ denotes the trainable parameters in the timestep embedding circuit. The architecture of an timestep embedding circuit with $N_{\text{time}}=4$ qubits is shown in Fig. 4. The circuit includes single-qubit gates $Rx$ and $Ry$, and the two-qubit gate $ZZ(\phi)=\exp{(-i\phi(Z\otimes Z)/2)}$. Our proposed method maps temporal information to appropriate positions on the Bloch sphere (as demonstrated by our numerical simulations) while maintaining linear complexity in the number of quantum gates, which is sufficient for our task requirements.
Figure 4: The architecture of the timestep embedding circuit $\mathcal{T}(\boldsymbol{\omega},t^{\prime})$ with trainable parameter $\boldsymbol{\omega}$, and given parameter $t^{\prime}=t\pi/T$. The circuit in the dashed lines is repeated $L_{\mathcal{T}}$ times. $l\in\{1,\dots,L_{\mathcal{T}}\}$.
III-C2 Denoising Circuit
@@
The denoising circuit $U(\boldsymbol{\theta})$ acts on the composite system $\tau_{t}\otimes\rho_{t}$. Subsequently, all qubits are regrouped into subsystems A and B, where they possess $N$ and $N_{\text{time}}$ qubits, respectively. By tracing out subsystem B, we obtain the predicted output $\tilde{\rho}_{t-1}$:
@@
Using register B to output ${\tilde{\rho}}_{t-1}$ might seem intuitive. However, such a design would result in the model failing. This failure is caused by the slight difference between $\rho_{t}$ and the ground truth $\rho_{t-1}$, which poses a training challenge. Rather than producing the intended output, the optimizer often merely copies its input. Section V provides a detailed examination of this issue, including numerical simulations to demonstrate the concept.
The architecture of a 4-qubit denoising circuit $U(\boldsymbol{\theta})$ is shown in Fig. 5. The circuit includes single-qubit gates $Rz$ and $Rx$, and the two-qubit gate $XX(\phi)=\exp{(-i\phi(X\otimes X)/2)}$. This circuit architecture using $\mathcal{O}(N^{2})$ two-qubit gates may affect the efficiency of our method. However, our focus is to evaluate the effectiveness of our method. To ensure that this evaluation is not affected by the circuit design, we choose a more complex architecture, even though it increases the gate complexity. This decision allows our results to accurately represent the potential of our method. In addition, it may not be necessary for us to maintain a consistent circuit architecture across all timesteps. Adaptable circuits that change architecture at different timesteps can reduce the total number of gates in QGDM. We leave the task of designing efficient circuit architectures for QGDM for future work. An effective approach is to use quantum architecture search algorithms [64, 65, 66, 67, 68, 69].
Figure 5: The architecture of a 4-qubit denoising circuit $U(\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ denotes the trainable parameters. To enhance the learning capability of the circuit, the part enclosed by the dashed lines is repeated $L$ times. $l\in\{1,\dots,L\}$.
III-D Training and Generation
@@
$\min_{\boldsymbol{\Theta}}\mathcal{L}=\mathcal{L}_{0}+\lambda\mathbb{E}_{t\in\text{Uniform}(\{2,\dots,T\})}\left[\mathcal{L}_{t-1}\right],$ (8)
where $\lambda$ is a hyperparameter used to balance the losses $\mathcal{L}_{0}$ and $\mathbb{E}_{t\in\text{Uniform}(\{2,\dots,T\})}\left[\mathcal{L}_{t-1}\right]$, and we find that setting a small $\lambda$ improves the generative effect of QGDM. $\text{Uniform}(\{2,\dots,T\})$ represents a uniform distribution that gives scalar values from 2 to $T$. $\mathcal{L}_{t-1}$ is calculated by:
@@
where $F(\cdot,\cdot)$ is the quantum fidelity function [35, 56], used to measure the distance between any two quantum states $\rho$ and $\sigma$:
@@
The QGDM training algorithm is presented in Algorithm 1. In each iteration, we compute the $\mathcal{L}_{0}$ and the expectation $\mathbb{E}_{t\in\text{Uniform}({2,\dots,T})}[\mathcal{L}_{t-1}]$ to update the trainable parameters until the training loss converges or the maximum number of iterations is reached. The expected value is estimated using the Monte Carlo approach.
Once the parameters are trained, we can iteratively obtain the target state $\rho_{0}$ from the completely mixed state $\rho_{T}$ using $\boldsymbol{\omega}$ and $\boldsymbol{\theta}$. The target quantum state $\rho_{0}$ is then generated using Algorithm 2. The algorithm starts with a completely mixed state and uses the trained denoising process to generate the target state step by step.
Algorithm 1 QGDM training algorithm.
1:  Input: Noise schedule parameters $\{\bar{\alpha}_{t}\}_{t=1}^{T}$, batch size BS, weight parameter $\lambda$, number of qubits of the target state $N$, number of qubits of the timestep embedding state $N_{\text{time}}$, total timestep of the diffusion process $T$, maximum iteration number, target quantum state $\rho_{0}$, learning rate $\eta$, completely mixed state $\mathbb{I}/d$.
2:  $\boldsymbol{\theta}\sim\text{Uniform}(0,\pi)$, $\boldsymbol{\omega}\sim\text{Uniform}(0,\pi)$
@@
5:     $t^{\prime}\leftarrow\pi/T$
6:     $\tau_{1}\leftarrow\mathcal{T}(\boldsymbol{\omega},t^{\prime})\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\text{time}}}\mathcal{T}^{\dagger}(\boldsymbol{\omega},t^{\prime})$
7:     $\tilde{\rho}_{0}\leftarrow\text{tr}_{\text{{B}}}\left(U(\boldsymbol{\theta})\left(\tau_{1}\otimes\rho_{1}\right)U^{\dagger}(\boldsymbol{\theta})\right)$
8:     $\mathcal{L}_{0}\leftarrow 1-F(\rho_{0},\tilde{\rho}_{0})$
9:     $\mathcal{L}\leftarrow 0$
10:     for $i$ from 1 to BS do
11:        Sample a unique timestep $t$ from the uniform distribution $\text{Uniform}(\{2,\dots,T\})$
12:        $\rho_{t}\leftarrow\left(1-\bar{\alpha}_{\mathrm{t}}\right)\mathbb{I}/d+\bar{\alpha}_{\mathrm{t}}\rho_{0}$
13:        $t^{\prime}\leftarrow t\pi/T$
14:        $\tau_{t}\leftarrow\mathcal{T}(\boldsymbol{\omega},t^{\prime})\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\text{time}}}\mathcal{T}^{\dagger}(\boldsymbol{\omega},t^{\prime})$
15:        $\tilde{\rho}_{t-1}\leftarrow\text{tr}_{\text{{B}}}\left(U(\boldsymbol{\theta})\left(\tau_{t}\otimes\rho_{t}\right)U^{\dagger}(\boldsymbol{\theta})\right)$
16:        $\rho_{t-1}\leftarrow\left(1-\bar{\alpha}_{\mathrm{t-1}}\right)\mathbb{I}/d+\bar{\alpha}_{\mathrm{t-1}}\rho_{0}$
17:        $\mathcal{L}\leftarrow\mathcal{L}+1-F\left(\rho_{t-1},\tilde{\rho}_{t-1}\right)$
18:     end for
19:     $\mathcal{L}\leftarrow\mathcal{L}_{0}+\lambda\frac{1}{\text{BS}}\mathcal{L}$
20:     $\boldsymbol{\omega}\leftarrow\boldsymbol{\omega}-\eta\nabla_{\boldsymbol{\omega}}\mathcal{L}$
21:     $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}$
22:  until $\mathcal{L}$ converges or the number of iterations reaches the maximum
23:  Output: Optimal parameters $\boldsymbol{\omega}^{*},\boldsymbol{\theta}^{*}$
Algorithm 2 QGDM generation algorithm.
1:  Input: Number of qubits of the timesetp embedding state $N_{\text{time}}$, total timestep of the diffusion process $T$, the optimal parameters $\boldsymbol{\omega}^{*},\boldsymbol{\theta}^{*}$, completely mixed state $\rho_{T}=\mathbb{I}/d$.
@@
5:     $\tau_{t}\leftarrow\mathcal{T}(\boldsymbol{\omega}^{*},t^{\prime})\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\text{time}}}\mathcal{T}^{\dagger}(\boldsymbol{\omega}^{*},t^{\prime})$
@@
IV Resource-Efficient Quantum Generative Diffusion Model (RE-QGDM)
The number of qubits required for the QGDM denoising process is $N_{\text{time}}+N$. To reduce the number of auxiliary qubits, we propose a new design for the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$. We call this prototype as the Resource-Efficient Quantum Generative Diffusion Model (RE-QGDM).
Figure 6: The denoising process design in Resource-Efficient Quantum Generative Diffusion Models (RE-QGDM).
In RE-QGDM, only $N+1$ qubits are needed to generate an $N$-qubit quantum state, which enhances its efficiency in terms of resource utilization. Fig. 6 illustrates the denoising process circuit for a 3-qubit target state using RE-QGDM. The polluted state $\rho_{t}$ is compressed by a parameterized circuit $U_{1}(\boldsymbol{\theta}_{1})$, with $\boldsymbol{\theta}_{1}$ as its trainable parameters, condensing its information into the final qubit. This qubit, containing concentrated information, is then combined with $\tau_{t}$ to form a composite system. Another circuit, $U_{2}(\boldsymbol{\theta}_{2})$, processes this expanded state across $N+1$ qubits. After applying $U_{2}(\boldsymbol{\theta}_{2})$, the system’s last qubit is discarded, leaving the reduced density matrix of the remaining $N$ qubits as the predicted output state ${\tilde{\rho}}_{{\rm{t}}-1}$:

### v3

@@ 3.2 Forward Process
In the natural world, diffusion is governed by non-equilibrium thermodynamics, driving systems toward equilibrium [56]. As a system undergoes evolution, its entropy increases until it reaches the maximum at equilibrium, indicating that there is no further evolution. The classical diffusion model draws inspiration from this diffusion mechanism [31]. In quantum information theory, the depolarizing channel is considered as a “worst-case scenario” [57]. Under its influence, a quantum state gradually loses its original information, eventually transforming into a completely mixed state. This state is characterized by the maximum entropy, representing the pinnacle of randomness in quantum systems. At this point, our uncertainty about the system has also reached its maximum.
@@
where $d=2^{N}$ denotes the dimension of Hilbert space and $\alpha_{t}\in[0,1]$ is a timestep-dependent scalar. Eq. (1) indicates that the input quantum state $\rho_{t-1}$ remains unchanged with a probability of $\alpha_{t}$, while with a probability of $1-\alpha_{t}$, it is replaced by the completely mixed state $\mathbb{I}/d$. This indicates the loss of original information in the quantum system. Therefore, the depolarizing channel is non-unitary because information is lost when quantum states pass through it, making the process irreversible. By setting an appropriate sequence of $\{\alpha_{t}\}$, a series of depolarizing channel sequences acting on the quantum state $\rho_{0}$ can eventually yield the output $\rho_{T}\approx\mathbb{I}/d$.
3.2.1 Efficient Quantum State Transition to Arbitrary Timestep
@@
where $\bar{\alpha}_{t}=\prod\limits_{i=1}^{t}\alpha_{i}$. Thus, $\bar{\alpha}_{t}\leq\bar{\alpha}_{t-1}$ since $\alpha_{i}\in[0,1],\forall i$.
@@
$=\left(1-\bar{\alpha}_{t}\right)\mathbb{I}/d+\bar{\alpha}_{t}\rho_{0}.~{}\square$
@@
3.2.2 Noise Schedule
The choice of ${\alpha_{t}}$ must ensure that $\bar{\alpha}_{T}\to 0$ and ${\alpha_{t}}\in\left[{0,1}\right]$ for all $t$. We use the cosine noise schedule from the classical diffusion model [34] to select $\alpha_{t}$:
@@ 3.2.2 Noise Schedule
with $s$ being a small offset hyperparameter. We note that there are other noise schedules, such as the linear noise schedule [32] and learned one [58]. Exploring the impacts of other noise schedules on QGDM’s performance, or designing specific noise schedules for QGDM, is an interesting research direction, which we leave for future work.
3.3 Backward Process
@@
The objective of a backward process in QGDM is to restore $\rho_{0}$ from $\rho_{T}$. We should first address two considerations.
@@
Firstly, since the denoising process is essentially the inverse of a non-unitary diffusion process, it must also be non-unitary to effectively denoise. Secondly, in principle, we need to train $T$ distinct denoising processes to deal with $T$ different denoising demands if we model the denoising process in a form of $f_{\boldsymbol{\Theta}_{t}}(\rho_{t})$. To reduce the number of trainable parameters in QGDM, we make these processes share the same parameters by using the timestep $t$ as an input. This integration is crucial for shaping $f_{\boldsymbol{\Theta}}(\rho_{t},t)$, as it allows the model to effectively utilize temporal information to accurately estimate $\rho_{t-1}$ while maintaining a small number of parameters.
@@
Figure 3: The denoising process framework, $f_{\boldsymbol{\Theta}}(\rho_{t},t)$, is composed of a timestep embedding circuit $\mathcal{T}(\boldsymbol{\omega},t^{\prime})$ and a denoising circuit $U(\boldsymbol{\theta})$, where $\boldsymbol{\Theta}=\{\boldsymbol{\omega},\boldsymbol{\theta}\}$.
@@
To endow the denoising process with the aforementioned two characteristics, we achieve non-unitary properties by tracing out a part of the subsystem. Additionally, we use a parameterized circuit to embed temporal information into the quantum system. Our design for the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ is illustrated in Fig. 3. $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ comprises two modules: a timestep embedding circuit $\mathcal{T}(\boldsymbol{\omega},t^{\prime})$ parameterized by $\boldsymbol{\omega}$ and $t^{\prime}=t\pi/T$, and a denoising circuit $U(\boldsymbol{\theta})$ parameterized by $\boldsymbol{\theta}$, thus trainable parameters $\boldsymbol{\Theta}=\{\boldsymbol{\omega},\boldsymbol{\theta}\}$. A timestep embedding circuit $\mathcal{T}(\boldsymbol{\omega},t^{\prime})$, operating on an auxiliary register with $N_{\tau}$ qubits, is utilized to produce state $\tau_{t}$, which carries temporal information. Subsequently, the composite system $\tau_{t}\otimes\rho_{t}$ is fed into the denoising circuit $U({\boldsymbol{\theta}})$. Afterward, all qubits are regrouped into two registers: A with $N$ qubits and B with $N_{\tau}$ qubits. We trace out B to get the predicted denoised quantum state $\tilde{\rho}_{t-1}$.
3.3.1 Timestep Embedding Circuit
@@
A critical consideration when designing the denoising process is how to encode the temporal information into a quantum state. This is a classical-to-quantum encoding challenge, where such methods as qubit encoding [59, 60] or amplitude encoding [61] could be used to embed $t$ into a quantum system. However, the qubit encoding method only maps the timestep sequence uniformly onto a ring on the Bloch sphere, which lacks expressiveness and flexibility. The amplitude encoding method requires a number of gates that is exponential in the number of qubits [62]. To enhance the expressiveness and flexibility of timestep embedding while maintaining an appropriate gate complexity, this work utilizes the quantum embedding method [63] to obtain a timestep embedding state $\tau_{t}$ with $N_{\tau}$ qubits, $i.e.$,
@@
$\tau_{t}=\mathcal{T}(\boldsymbol{\omega},t^{\prime})\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\tau}}\mathcal{T}^{\dagger}(\boldsymbol{\omega},t^{\prime}),$ (6)
where scalar $t^{\prime}=t\pi/T$ is the mapping of timestep $t$ into the range [0, $\pi$], and $\boldsymbol{\omega}$ denotes the trainable parameters in the timestep embedding circuit.
The architecture of the timestep embedding circuit is shown in the supplementary file.
3.3.2 Denoising Circuit
@@
The denoising circuit $U(\boldsymbol{\theta})$ acts on the composite system $\tau_{t}\otimes\rho_{t}$. Subsequently, all qubits are regrouped into subsystems A and B, where they possess $N$ and $N_{\tau}$ qubits, respectively. By tracing out B, we obtain the predicted output:
@@
Using subsystem B to output ${\tilde{\rho}}_{t-1}$ seems intuitive. However, such a design would result in the model failure. This failure is caused by the slight difference between $\rho_{t}$ and ground truth $\rho_{t-1}$, which poses a training challenge. Rather than producing the intended output, the optimizer often merely copies its input. The next Section provides the detailed examination of this issue, including numerical simulations to demonstrate the concept.
For details on the architecture of the denoising circuit, please refer to the supplementary file.
3.4 Training and Generation
@@
$\min_{\boldsymbol{\Theta}}\mathcal{L}=\mathcal{L}_{0}+\lambda\mathbb{E}_{t\in\mathcal{U}(2,\dots,T)}\left[\mathcal{L}_{t-1}\right],$ (8)
where $\lambda$ is a hyperparameter used to balance the losses $\mathcal{L}_{0}$ and $\mathbb{E}_{t\in\mathcal{U}(2,\dots,T)}\left[\mathcal{L}_{t-1}\right]$, and we find that setting a small $\lambda$ improves the generative effect of QGDM. $\mathcal{U}(2,\dots,T)$ represents a uniform distribution that gives values from 2 to $T$. We have:
@@
where $F(\cdot,\cdot)$ is the quantum fidelity function [38, 57], used to measure the distance between two quantum states $\rho$ and $\sigma$:
@@
QGDM training is realized via in Algorithm 1. In each iteration, we compute $\mathcal{L}_{0}$ and the expectation $\mathbb{E}_{t\in\mathcal{U}(2,\dots,T)}[\mathcal{L}_{t-1}]$ to update the trainable parameters until the training loss converges or the maximum number of iterations is reached. The expected value is estimated by using the Monte Carlo approach.
Once the parameters are trained, we can iteratively obtain the target state $\rho_{0}$ from the completely mixed state $\rho_{T}$ by using $\boldsymbol{\omega}$ and $\boldsymbol{\theta}$. $\rho_{0}$ is then generated via Algorithm 2. The algorithm starts with a completely mixed state and uses the trained denoising process to generate the target state via $T$ steps. Therefore, its complexity is $\mathcal{O}(T)$.
Algorithm 1 QGDM training.
1:  Input: Noise schedule parameters $\{\bar{\alpha}_{t}\}_{t=1}^{T}$, batch size BS, weight parameter $\lambda$, number of qubits of the target state $N$, number of qubits of the timestep embedding state $N_{\tau}$, total timestep of the diffusion process $T$, maximum iteration number, target quantum state $\rho_{0}$, learning rate $\eta$, completely mixed state $\mathbb{I}/d$.
2:  $\boldsymbol{\theta}\sim\mathcal{U}(0,\pi)$, $\boldsymbol{\omega}\sim\mathcal{U}(0,\pi)$
@@
5:     $\tau_{1}\leftarrow\mathcal{T}(\boldsymbol{\omega},\pi/T)\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\tau}}\mathcal{T}^{\dagger}(\boldsymbol{\omega},\pi/T)$
6:     $\tilde{\rho}_{0}\leftarrow\text{tr}_{\text{{B}}}\left(U(\boldsymbol{\theta})\left(\tau_{1}\otimes\rho_{1}\right)U^{\dagger}(\boldsymbol{\theta})\right)$
7:     $\mathcal{L}_{0}\leftarrow 1-F(\rho_{0},\tilde{\rho}_{0})$
8:     $\mathcal{L}\leftarrow 0$
9:     for $i$ from 1 to BS do
10:        Sample a unique timestep $t$ from the uniform distribution $\mathcal{U}(2,\dots,T)$
11:        $\rho_{t}\leftarrow\left(1-\bar{\alpha}_{\mathrm{t}}\right)\mathbb{I}/d+\bar{\alpha}_{\mathrm{t}}\rho_{0}$
12:        $\tau_{t}\leftarrow\mathcal{T}(\boldsymbol{\omega},t\pi/T)\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\tau}}\mathcal{T}^{\dagger}(\boldsymbol{\omega},t\pi/T)$
13:        $\tilde{\rho}_{t-1}\leftarrow\text{tr}_{\text{{B}}}\left(U(\boldsymbol{\theta})\left(\tau_{t}\otimes\rho_{t}\right)U^{\dagger}(\boldsymbol{\theta})\right)$
14:        $\rho_{t-1}\leftarrow\left(1-\bar{\alpha}_{\mathrm{t-1}}\right)\mathbb{I}/d+\bar{\alpha}_{\mathrm{t-1}}\rho_{0}$
15:        $\mathcal{L}\leftarrow\mathcal{L}+1-F\left(\rho_{t-1},\tilde{\rho}_{t-1}\right)$
16:     end for
17:     $\mathcal{L}\leftarrow\mathcal{L}_{0}+\lambda\frac{1}{\text{BS}}\mathcal{L}$
18:     $\boldsymbol{\omega}\leftarrow\boldsymbol{\omega}-\eta\nabla_{\boldsymbol{\omega}}\mathcal{L};\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}$
19:  until $\mathcal{L}$ converges or the number of iterations reaches the maximum
20:  Output: Optimal parameters $\boldsymbol{\omega}^{*},\boldsymbol{\theta}^{*}$
Algorithm 2 QGDM generation.
1:  Input: Number of qubits of the timesetp embedding state $N_{\tau}$, total timestep of the diffusion process $T$, the optimal parameters $\boldsymbol{\omega}^{*},\boldsymbol{\theta}^{*}$, completely mixed state $\rho_{T}=\mathbb{I}/d$.
@@
5:     $\tau_{t}\leftarrow\mathcal{T}(\boldsymbol{\omega}^{*},t^{\prime})\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\tau}}\mathcal{T}^{\dagger}(\boldsymbol{\omega}^{*},t^{\prime})$
@@
4 Resource-efficient Quantum Generative Diffusion Model (RQGDM)
The number of qubits required for the QGDM denoising process is $N_{\tau}+N$. To reduce the number of auxiliary qubits, we propose a new design for the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$. We call this prototype as RQGDM.
Figure 4: The backward process design in Resource-efficient Quantum Generative Diffusion Models (RQGDM).
Inspired by quantum data compression methods [64, 65], the core idea of RQGDM is to first use a parameterized quantum circuit to learn a low-dimensional representation of quantum data from its original high-dimensional Hilbert space for the denoising process. This low-dimensional representation uses fewer qubits than the input state (e.g., one qubit in our approach), thereby reducing the total number of qubits required for the denoising process. In RQGDM, only $N+1$ qubits are needed to generate an $N$-qubit quantum state, which enhances its efficiency in terms of resource utilization. Fig. 4 illustrates the denoising process circuit for a 3-qubit target state using RQGDM. The polluted state $\rho_{t}$ is compressed by a parameterized circuit $U_{1}(\boldsymbol{\theta}_{1})$, with $\boldsymbol{\theta}_{1}$ as its trainable parameters, condensing its information into the final qubit. This qubit, containing concentrated information, is then combined with $\tau_{t}$ to form a composite system. Another circuit, $U_{2}(\boldsymbol{\theta}_{2})$, processes this expanded state across $N+1$ qubits. After applying $U_{2}(\boldsymbol{\theta}_{2})$, the system’s last qubit is discarded, leaving the reduced density matrix of the remaining $N$ qubits as the predicted output state ${\tilde{\rho}}_{{\rm{t}}-1}$: